OpenStudy (anonymous):

please help:) What is the simplified form of (x^2yx)^2(xy^2z^2) (xy^2) A. (xyz)^2(xyz)/x B. 1/(xyz)^2 C. x^3y^2z^2 D. x^4y^4z

OpenStudy (callisto):

Is it \[\frac{(x^2yx)^2(xy^2z^2)}{(xy^2)}\] for your question?

OpenStudy (anonymous):

yeah

OpenStudy (callisto):

There are 3 things you need to know for this question 1. \[(x^m)^n = x^{mn}\] 2. \[x^mx^n = x^{m+n}\] 3. \[\frac{x^m}{x^n}=x^{m-n}\]

OpenStudy (anonymous):

i really just need the answer cuz i cant figure this one out for my homework

OpenStudy (diyadiya):

Please Read http://openstudy.com/code-of-conduct

OpenStudy (callisto):